OpenStudy (anonymous):

Find the value of 3 over the square root of 3.. 3 3 times the square root of 3. the square root of 3. the square root of 3 over 3.

OpenStudy (abb0t):

\(\frac{3}{\sqrt{3}}\)?

OpenStudy (anonymous):

\[\frac{3}{\sqrt3} = \frac{3}{\sqrt3} \times \frac{\sqrt 3}{\sqrt3} =\frac{3\sqrt 3}{\sqrt 3 \times \sqrt3} \] \[=\frac{3\sqrt 3}{3} =\sqrt 3\] Hence the square root of 3. is the correct Ans